### 17 Jan 2024 v4

@@ A.1 The natural thin plate spline smoother
The natural thin plate spline problem provides a smooth function $g$ such that $g(\boldsymbol{\xi}_{i})\approx y_{i}$ for $i=1,\dotsc,n$. The “closeness” of $g(\boldsymbol{\xi}_{i})$ to $y_{i}$ is enforced through squared error, while a penalty on the curvature of $g$ discourages roughness. Formally, we define the roughness penalty:
@@
where $\boldsymbol{\delta}\in\mathbb{R}^{n},\boldsymbol{\beta}\in\mathbb{R}^{3}$ uniquely define the thin plate spline $g$. Expanding (35), we obtain:
@@
A.2 The covariance-reweighed thin plate spline smoother
Due to the correlated and heteroskedastic nature of the temperature differences in this paper, we desire a version of the thin plate spline smoother (37) that properly accounts for the these effects. This is done via generalized least squares (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003), which poses the objective function:
@@ A.2 The covariance-reweighed thin plate spline smoother
where $\mathbf{W}=\operatorname{\mathrm{cov}}(\boldsymbol{y})^{-1}$ is the inverse of the covariance matrix between the seasonally adjusted temperature differences, previously estimated in Section 3.4. In the simplest case of diagonal $\mathbf{W}$, (38) may be interpreted as inversely weighing the contribution of each observation to the objective by the magnitude of its variance.
@@ A.3 Adaptation to fixed knots
Finally, we detail the adaptation of the covariance-reweighed thin plate spline smoother to the fixed-knot regime (Wahba, 1980, 1990; Ruppert et al., 2003; Wood, 2017). Suppose a set of fixed knots $\tilde{\boldsymbol{\xi}}_{1},\dotsc,\tilde{\boldsymbol{\xi}}_{m}$, with $\tilde{\mathbf{R}}$ as defined in (23) and $\mathbf{S}$} as defined in (25). The thin plate spline smoothing problem, restricted to this set of knots, is given (Wood, 2017) by:
@@ Appendix B Leave-one-out cross-validation
where $w_{i}$ is an observation-specific weight (typically the inverse variance) and $\hat{g}^{(-i)}(\boldsymbol{\xi}_{i};\lambda)$ is the thin plate spline, fitted to all observations except observation $i$, using regularization parameter $\lambda$, evaluated on observation $i$.
As the thin plate spline is a linear smoother, we may trivially access the leave-one-out residuals through the “hat matrix trick” (Cook and Weisberg, 1982; Green and Silverman, 1994)
@@
Figure 13: We observe Figure 13(a) that the minimizing $\hat{\lambda}^{\text{min}}$, for each depth, tends to lie in an extremely shallow basin, within which a wide range of $\lambda$ achieve comparable estimated test error (n.b., the vertical axis is on the logarithmic scale). By choosing a $\hat{\lambda}^{\text{1.01}}$ associated with a 1% increase in test error, we obtain fits that more faithfully illustrate the TC-induced thermal response. Figure 13(b) reports the chosen $\hat{\lambda}$ for each pressure level $z$ and for the vertically averaged temperature differences. Fits using $\lambda>500$ are ignored as they essentially only fit the linear (unpenalized) terms of the thin plate spline.

### 14 Mar 2024 v5

@@ A.1 The natural thin-plate spline smoother
The natural thin-plate spline problem provides a smooth function $g$ such that $g(\boldsymbol{\xi}_{i})\approx y_{i}$ for $i=1,\dotsc,n$. The “closeness” of $g(\boldsymbol{\xi}_{i})$ to $y_{i}$ is enforced through squared error, while a penalty on the curvature of $g$ discourages roughness. Formally, we define the roughness penalty:
@@
where $\boldsymbol{\delta}\in\mathbb{R}^{n},\boldsymbol{\beta}\in\mathbb{R}^{3}$ uniquely define the thin-plate spline $g$. Expanding (35), we obtain:
@@
A.2 The covariance-reweighted thin-plate spline smoother
Due to the correlated and heteroskedastic nature of the temperature differences in this paper, we desire a version of the thin-plate spline smoother (37) that properly accounts for the these effects. This is done via generalized least squares (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003), which poses the objective function:
@@ A.2 The covariance-reweighted thin-plate spline smoother
where $\mathbf{W}=\operatorname{\mathrm{cov}}(\boldsymbol{y})^{-1}$ is the inverse of the covariance matrix between the seasonally adjusted temperature differences, previously estimated in Section 3.4. In the simplest case of diagonal $\mathbf{W}$, (38) may be interpreted as inversely weighting the contribution of each observation to the objective by the magnitude of its variance.
@@ A.3 Adaptation to fixed knots
Finally, we detail the adaptation of the covariance-reweighted thin-plate spline smoother to the fixed-knot regime (Wahba, 1980, 1990; Ruppert et al., 2003; Wood, 2017). Suppose a set of fixed knots $\tilde{\boldsymbol{\xi}}_{1},\dotsc,\tilde{\boldsymbol{\xi}}_{m}$, with $\tilde{\mathbf{R}}$ as defined in (23) and $\mathbf{S}$} as defined in (25). The thin-plate spline smoothing problem, restricted to this set of knots, is given (Wood, 2017) by:
@@ Appendix B Leave-one-out cross-validation
where $w_{i}$ is an observation-specific weight (typically the inverse variance) and $\hat{g}^{(-i)}(\boldsymbol{\xi}_{i};\lambda)$ is the thin-plate spline, fitted to all observations except observation $i$, using regularization parameter $\lambda$, evaluated on observation $i$.
As the thin-plate spline is a linear smoother, we may trivially access the leave-one-out residuals through the “hat matrix trick” (Cook and Weisberg, 1982; Green and Silverman, 1994)
@@
Figure 13: We observe Figure 13(a) that the minimizing $\hat{\lambda}^{\text{min}}$, for each depth, tends to lie in an extremely shallow basin, within which a wide range of $\lambda$ achieve comparable estimated test error (n.b., the vertical axis is on the logarithmic scale). By choosing a $\hat{\lambda}^{\text{1.01}}$ associated with a 1% increase in test error, we obtain fits that more faithfully illustrate the TC-induced thermal response. Figure 13(b) reports the chosen $\hat{\lambda}$ for each pressure level $z$ and for the vertically averaged temperature differences. Fits using $\lambda>500$ are ignored as they essentially only fit the linear (unpenalized) terms of the thin-plate spline.